shonda:

Alex wants to start an IRA that will have $1,250,675 in it when he retires in 30 years. How much should he invest semiannually in his IRA to do this if the interest is 4.5% compounded semiannually? Assume an Annuity Due. Round to the nearest cent.

jimthompson5910:

https://www.financeformulas.net/Future-Value-of-Annuity-Due.html The formula we'll use is the future value of annuity due The formula is \[FV = (1+r)*P*\left[ \frac{(1+r)^n-1}{r}\right]\] Keep in mind that this r value is not the annual rate, and instead it is the semiannual rate (since the deposit periods are semiannual) An annual rate of 4.5% cuts in half to 2.25% so r = 0.0225 n = number of payment periods (ie number of deposits) n = 30*2 = 60 payments will be deposited P = unknown payment per period, which is what we want to solve for FV = 1,250,675 = amount we want in the account after 30 years So, \[FV = (1+r)*P*\left[ \frac{(1+r)^n-1}{r}\right]\] \[1250675 = (1+0.025)*P*\left[ \frac{(1+0.025)^{60}-1}{0.025}\right]\] I'll let you solve for P